# Leveraging Nested MLMC for Sequential Neural Posterior Estimation with Intractable Likelihoods

### v1

@@ 5 Numerical experiments
In this section, numerical experiments are conducted to compare the efficiency of two unbiased MLMC methods and a biased MLMC method for three models. To handle the curse of dimensionality, we propose to use low-dimensional summary statistics instead of the full data. However, this introduces additional bias, as detailed in [15]. In this paper, we do not look into the effects of using summary statistics for SNPE methods. We refer to Fearnhead and Prangle [15] for a semi-automatic method of constructing summary statistics in the context of ABC.
A toy example. Two-moon model was studied in Greenberg et al. [25]. For a given parameter $\theta\in\mathbb{R}^{2}$, the Two-moon simulator generates observations $x\in\mathbb{R}$ via
$a\sim U(-\frac{\pi}{2},\frac{\pi}{2})$ $,\quad r_{2}\sim\mathcal{N}(0.1,0.01^{2}),$
@@
The intermediate variables $p$ follow a single crescent-shaped distribution, which is then shifted and rotated around the origin based on the parameter values of $\theta$. The absolute value $|\theta_{1}+\theta_{2}|$ contributes to the emergence of a second crescent in the posterior distribution. We choose a uniform prior over the square $[-1,1]^{2}$ to perform the inference.
Lotka-Volterra model. This model describes the continuous time evolution of a population of predators interacting with a population of prey using a stochastic Markov jump process. The model describe that, the birth of a predator at a rate $\exp(\theta_{1})XY$, resulting in an increase of $X$ by one; The death of a predator at a rate proportional to $\exp(\theta_{2})X$, leading to a decrease of $X$ by one; The birth of a prey at a rate proportional to $\exp(\theta_{3})Y$, resulting in an increase of $Y$ by one; The consumption of a prey by a predator at a rate proportional to $\exp(\theta_{4})XY$, leading to a decrease of $Y$ by one. Following the experimental details outlined in [49], we initialize the predator and prey populations as $X=50$ and $Y=100$, respectively. We conduct simulations of the Lotka-Volterra model using the Gillespie algorithm [21] over a duration of 30 time units. We recorded the populations at intervals of 0.2 time units, resulting in time series data sets, each consisting of 151 values. The resulting summary statistics $S(x)$ are represented as a 9-dimensional vector, which includes the following time series features: the logarithm of the mean of each time series, the logarithm of the variance of each time series, the auto-correlation coefficient of each time series at lags of 0.2 and 0.4 time units, and the cross-correlation coefficient between the two time series. In our experiments, the prior distribution of the parameters is set to $\mathcal{U}(-5,2)^{4}$, and we generate the ground truth posterior with SMC-ABC [4], which is more costly than methods in this paper.
@@
In our experiments, we set $I=50$ jobs and the summary statistics $S(x)$ has been selected as the logarithm of 0th, 25th, 50th, 75th and 100th percentiles of the set of inter-departure times. The prior distribution of the parameters is
@@
In the training process, we simulate $N=1000$ samples in each round, and with $R=20$ rounds in total. In each round, we randomly pick 5% of the newly generated samples $\theta$ and their corresponding $x$ values as validation data. We follow the early stop criterion proposed by [49], which terminates the training if the loss value on the validation data does not decrease after 20 epochs in a single round. For the optimizer, we use Adam [36] with a batch size of 100, a learning rate of $1\times 10^{-4}$, and a weight decay of $1\times 10^{-4}$.
@@
In this paper, we employ neural spline flows (NSFs) [14] as the conditional density estimator, which consists of 8 layers. Each layer is constructed using two residual blocks with 50 units and ReLU activation function. With 10 bins in each monotonic piecewise rational-quadratic transform, and the tail bound is set to 20.
@@
We compare the performance of RU-MLMC, GRR-MLMC, and TGRR-MLMC on the Two-moon, Lotka-Volterra, and M/G/1 models. To assess the similarity between the approximate posterior distribution $q_{F(x_{o},\phi)}(\theta)$ and the true posterior distribution $p(\theta|x_{o})$ given the observed data, we employ maximum mean discrepancy (MMD) [25, 26, 31, 49] and classifier two-sample tests (C2ST) [12, 28, 40] as discriminant criteria. Additionally, we use log median distance (LMD) to measure the distance between $x_{o}$ and $x$ drawn from $p(x|\theta)$, where $\theta$ is sampled from $q_{F(x_{o},\phi)}(\theta)$. In cases where the true posterior distribution $p(\theta|x_{o})$ is intractable even with knowledge of the sample generation process, we resort to the negative log probability (NLOG) [14, 25, 31, 48, 49] of the true parameters $\theta^{*}$; in this scenario, the observation $x_{o}$ is sampled from $p(x|\theta^{*})$. It is important to note that lower values for all the mentioned indicators are favorable.
For the setting of MLMC methods, we have chosen $M_{0}=8$ and $\overline{m}=4$ for TGRR-MLMC. The base level for GRR and TGRR has been set to ${\underline{m}}=2$.
@@
We now discuss the choice of the hyperparameter $\alpha$ for the geometric distribution of the level $L$. Unlike previous work where $\alpha$ is chosen to minimize the average cost $\mathrm{Cost}^{\mathrm{RU}}$ [22, 23, 30], based on experiment results in Figure 2, in addition to average cost, variance is another crucial factor that demands our attention. These two indicators collectively hold great significance in our evaluation and optimization processes. In order to reduce the variance of this estimate, one direct approach is to take average of different iterations. The asymptotic inefficiency [35] defined as $H_{\mathrm{RU}}:=\mathrm{Var}[V^{\mathrm{RU}}]\times\mathrm{Cost}_{\mathrm{RU}}$ remains a constant whether iteration average is taken. It would be effective to decrease variance by raising cost if $\alpha$ is chosen to minimize this quantity. Utilizing Theorem 3.5, we find an upper bound for the asymptotic inefficiency
$H_{\mathrm{RU}}\lesssim\frac{M_{0}2^{\alpha+r_{2}}}{\left(2^{\alpha}-2\right)\left(2^{r_{2}}-2^{\alpha}\right)}.$ (29)
@@
As a result, the optimal $\alpha$ minimizing the upper bound is $\alpha_{\mathrm{RU}}^{*}=(r_{2}+1)/2\in(1,r_{2})$. When ${\underline{m}}=0$ in GRR-MLMC, which degenerates to RR-MLMC, the constants given in Theorem 3.6 do not involve the leverage of Theorem 3.4. Therefore, these constants are the same as those in Theorem 3.5, which yields an uniform upper bound for $H_{\mathrm{RR}}$. Therefore, we reach the same conclusion for $\alpha^{\mathrm{RR}}_{\mathrm{opt}}$. From this point of view, we can see that RR-MLMC actually trades lower variance for higher average cost compared with RU-MLMC. We apply the same procedure for other MLMC methods, the corresponding optimal $\alpha$ are presented in Table 1.
@@
Since the proposal distributions of each round are distinct, the hyperparameter $r_{2}$ in Theorems 3.3 and 3.2 differs in each round. To determine the value of $r_{2}$ for each round, except for the first two, one may take the value from the previous round. To address this issue, we have conducted 50 training processes and sequentially performed linear regression, a universal value is then selected as $r_{2}=1.8$.
A
B
C
Figure 3: Density plot for RU-MLMC, GRR-MLMC and TGRR-MLMC A. Two-moon, from left to right: available ground truth, RU-MLMC, GRR-MLMC and TGRR-MLMC. B. Lotka-Volterra, from left to right: ground truth simulated with SMC-ABC [4], RU-MLMC, GRR-MLMC and TGRR-MLMC. C. M/G/1 queue model, the setting is the same with Lotka-Volterra.
@@
We present our results in Figures 3 and 4. It is observed that both unbiased methods are inferior to the biased one in some cases. When comparing RU-MLMC with GRR-MLMC in the case of the Lotka-Volterra and M/G/1 queue models, we find that the unbiased method can greatly benefit from variance reduction. In the case of Lotka-Volterra, when comparing TGRR and GRR methods, we conclude that the gradient information $\Delta\rho_{\phi,\ell}$ where $\ell\geq\overline{m}$ not only does not contribute to the overall performance, but also make it worse when measured with C2ST. As suggested by Theorem 4.4, this could be due to the domination of variance, where the effect of reducing variance is superior to reducing bias.
In conclusion, in the case of SNPE where the complex model of $q_{F(x,\phi)}(\theta)$ is used for density estimation, variance tends to dominate the bias as it is suggested by Theorem 4.4, indicating that excessive variance could seriously affects the training process of the density estimator. Instead of seeking for unbiasedness, one should try to strike a balance between bias, average computational cost, and variance of the gradient in this case.
It is worth noting that in the case of the Two-moon problem, RU-MLMC method deliver comparable performance to the other two methods with lower computational cost and time requirements. Rather than employing any of these methods indiscriminately, it is crucial to undertake a comprehensive and thoughtful analysis of the specific problem at hand.
@@ 6 Concluding remarks
We develop a series of efficient nested MLMC methods, which are improvement over the unbiased MLMC methods. The latter suffer from excessive variance at the cost of unbiasedness, which, based on the convergence analysis, severely affects the method’s performance. Through experiments on standard benchmark for likelihood-free estimation and theoretical analysis, we valid the inferior performance of unbiased methods when variance dominates bias. In this paper, our choice for hyper-parameter $\alpha$ deviates from the mainstream choice. While values of $\alpha$ closer to $r_{2}$ is favored, aiming at cost minimization, given the fact that variance is the major focus, our focus lies on the asymptotic inefficiency by considering both variance and average cost. However, given that a thorough validation for this strategy through ablation experiments is still lacking, its improvement is still unclear. Quasi-Monte Carlo (QMC) and Randomized Quasi-Monte Carlo (RQMC) are well known variance reduction techniques, its application in nested MLMC methods has already been investigated in [30]. Given the high sensitivity of this problem to variance, using RQMC is likely to be beneficial, and existing results [30] can be appropriately utilized.
Acknowledgments
The work was supported by the National Natural Science Foundation of China grant 12071154, the Guangdong Basic and Applied Basic Research Foundation grant 2021A1515010275.
Appendix A Supplementary proofs
A.1 Proof for Theorem 3.3
Theorem A.1.
Assume that there exists $s>2$ that
@@ Appendix A Supplementary proofs
$\underset{x,\phi,\theta}{\sup}\|\nabla_{\phi}\log g_{\phi}(x,\theta)\|_{\infty}=M_{max}<\infty,\,\underset{\phi}{\sup}\ \mathbb{E}\left[\left|\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right|^{s}\right]<\infty,$
then we have
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\right]$ $\leq\frac{16(2^{r_{2}}+1)(E_{1}+E_{2})H(2r)+9E_{3}H(s)1\{s<4\}}{(M_{0}2^{\ell})^{r_{2}}},$ (30)
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|\right]$ $\leq\frac{(6\sqrt{2}d^{1/4}M^{1/2}_{\max}+4d^{1/2}M_{\max})C_{2}H(2)}{M_{\ell}},$ (31)

### v2

@@ 5 Numerical experiments
In this section, numerical experiments compare the efficiency of two unbiased MLMC methods and a biased MLMC method for three models. To handle high-dimensional datasets, it is common to use informative low-dimensional summary statistics of the datasets instead. However, this introduces additional bias, as detailed in [14]. In this paper, we do not look into the effects of using summary statistics for SNPE methods. We refer to Fearnhead and Prangle [14] for a semi-automatic method of constructing summary statistics in the context of ABC.
A toy example. Two-moon model was studied in Greenberg et al. [24]. For a given parameter $\theta\in\mathbb{R}^{2}$, the Two-moon simulator generates observations $x\in\mathbb{R}^{2}$ via
$a\sim\mathcal{U}(-\pi/2,\pi/2)$ $,\quad r_{2}\sim\mathcal{N}(0.1,0.01^{2}),$
@@
The intermediate variables $p$ follow a single crescent-shaped distribution, which is then shifted and rotated around the origin based on the parameter values of $\theta$. The absolute value $|\theta_{1}+\theta_{2}|$ contributes to the emergence of a second crescent in the posterior distribution. We choose a uniform prior over the square $[-1,1]^{2}$ to perform the inference. The observation for this task is $x_{o}=(0,0)$ following [24], the ground truth parameters are $\theta^{*}=(0.2475,0.2475).$
Lotka-Volterra model. This model describes the continuous time evolution of a population of predators interacting with a population of prey using a stochastic Markov jump process. The model describes that the birth of a predator at a rate $\exp(\theta_{1})XY$, resulting in an increase of $X$ by one. The death of a predator at a rate proportional to $\exp(\theta_{2})X$, leading to a decrease of $X$ by one. The birth of a prey at a rate proportional to $\exp(\theta_{3})Y$, resulting in an increase of $Y$ by one. The consumption of a prey by a predator at a rate proportional to $\exp(\theta_{4})XY$, leading to a decrease of $Y$ by one. Following the experimental details outlined in [48], we initialize the predator and prey populations as $X=50$ and $Y=100$, respectively. We conduct simulations of the Lotka-Volterra model using Gillespie’s algorithm [20] throughout 30 time units. We recorded the populations at intervals of 0.2 time units, resulting in time series data sets, each consisting of 151 values. The resulting summary statistics $S(x)$ are represented as a 9-dimensional vector, which includes the following time series features: the logarithm of the mean of each time series, the logarithm of the variance of each time series, the auto-correlation coefficient of each time series at lags of 0.2 and 0.4-time units, and the cross-correlation coefficient between the two time series. In our experiments, the prior distribution of the parameters is set to $\mathcal{U}(-5,2)^{4}$, and we generate the ground truth posterior with SMC-ABC [4], which is more costly than methods in this paper. In our experiments, the prior distribution of the parameters is set to $\mathcal{U}(-5,2)^{4}$, the ground truth parameters are
$\theta^{*}=(\log 0.01,\ \log 0.5,\ \log 1,\ \log 0.01),$
the observed summary statistics $S(x_{o})$ simulated from the model with the ground truth parameters $\theta^{*}$ are
[…]
and the corresponding standard deviation under the ground truth parameters $\theta^{*}$ (based on 10,000 simulations) are
[…]
@@
In our experiments, we set $I=50$ jobs and the summary statistics $S(x)$ has been selected as the logarithm of 0th, 25th, 50th, 75th and 100th percentiles of the set of inter-departure times. The prior distribution of the parameters is given by
@@
A
B
C
Figure 2: Approximated posterior for RU-MLMC, GRR-MLMC and TGRR-MLMC A. Two-moon, from left to right: available ground truth, RU-MLMC, GRR-MLMC and TGRR-MLMC. B. Lotka-Volterra, from left to right: ground truth simulated with SMC-ABC [4], RU-MLMC, GRR-MLMC and TGRR-MLMC. C. M/G/1 queue model, the setting is the same with Lotka-Volterra. We have plotted the scatter plot for each of the 2D sub-spaces. The histogram for each $\theta_{i},i=1,\dots,4$ is plotted on the diagonal with the ground truth parameter marked with dotted lines.
@@
In the training process, we simulate $N=1,000$ samples in each round, and with $R=20$ rounds in total. In each round, we randomly pick 5% of the newly generated samples $\theta$ and their corresponding $x$ values as validation data. We follow the early stop criterion proposed by [48], which terminates the training if the loss value on the validation data does not decrease after 20 epochs in a single round. For the optimizer, we use Adam [35] with a batch size of 100, a learning rate of $1\times 10^{-4}$, and a weight decay of $1\times 10^{-4}$. The detailed track of the computational time of various methods can be found in Table 3.
@@
In this paper, we employ neural spline flows (NSFs) [13] as the conditional density estimator, which consists of 8 layers. Each layer is constructed using two residual blocks with 50 units and ReLU activation function. With 10 bins in each monotonic piecewise rational-quadratic transform the tail bound is set to 20. For a detailed introduction to the structure of NSFs, we refer the reader to Appendix B.
@@
We compare the performance of RU-MLMC, GRR-MLMC, and TGRR-MLMC on the Two-moon, Lotka-Volterra, M/G/1 models. To assess the similarity between the approximate posterior distribution $q_{F(x_{o},\phi)}(\theta)$ and the true posterior distribution $p(\theta|x_{o})$ given the observed data, we employ maximum mean discrepancy (MMD) [24, 25, 30, 48] and classifier two-sample tests (C2ST) [11, 27, 39] as discriminant criteria. Furthermore, we use the logarithmic median distance (LMD) to measure the distance between $x_{o}$ and $x$ drawn from $p(x|\theta)$, where $\theta$ is sampled from $q_{F(x_{o},\phi)}(\theta)$. The negative log-density (NLOG) value of the ground truth parameter $\theta^{*}$ of the generated observation $x_{o}$ under the conditional density structure, i.e., $-\log q_{F(x_{o},\phi)}(\theta^{*})$, is another useful metric. A lower value indicates that, in the parameter space, the posterior density estimator has a higher probability density at the true parameter, which can be used to evaluate the model’s accuracy in estimating the true parameter given the observed data [13, 24, 30, 47, 48].
For the setting of MLMC methods, we take $M_{0}=8$ and $\overline{m}=4$ for TGRR-MLMC. The base level for GRR and TGRR is ${\underline{m}}=2$. Since the proposal distributions of each round are distinct, the hyperparameter $r_{2}$ in Theorem 3.2 differs in each round. To determine the value of $r_{2}$ for each round, except for the first two, one may take the value from the previous round. To address this issue, we have conducted 50 training processes and sequentially performed linear regression, a universal value is then selected as $r_{2}=1.8$.
@@
We now discuss the choice of the hyperparameter $\alpha$ for the geometric distribution of the level $L$. Unlike previous work where $\alpha$ is chosen to minimize the average cost $\mathrm{Cost}_{\mathrm{RU}}$ [21, 22, 29], based on experiment results in Figure 1, in addition to average cost, variance is another crucial factor that demands our attention. These two indicators collectively hold great significance in our evaluation and optimization processes. We now try to minimize the asymptotic inefficiency [34] defined as $H_{\mathrm{RU}}:=\mathrm{Var}[V^{\mathrm{RU}}]\times\mathrm{Cost}_{\mathrm{RU}}$. Utilizing Theorem 3.3, we find an upper bound for the asymptotic inefficiency
$H_{\mathrm{RU}}\leq\frac{B2^{\alpha+r_{2}}}{\left(2^{\alpha}-2\right)\left(2^{r_{2}}-2^{\alpha}\right)}$ (34)
@@
for some constant $B>0$. As a result, the optimal $\alpha$ minimizing the upper bound is $\alpha_{\mathrm{RU}}^{*}=(r_{2}+1)/2\in(1,r_{2})$. When ${\underline{m}}=0$ in GRR-MLMC, which degenerates to RR-MLMC, we reach the same conclusion for $\alpha^{*}_{\mathrm{RR}}$. From this point of view, we can see that RR-MLMC actually trades lower variance for higher average cost compared with RU-MLMC. We apply the same procedure for other MLMC methods, and the corresponding optimal $\alpha$ are presented in Table 1. For an ablation study on the choices of $\alpha$, we refer the reader to Appendix D.
@@
We present our results for the baseline models in Figure 2 and Figure 3. It is observed that both unbiased methods are inferior to the biased ones in some cases. When comparing RU-MLMC with GRR-MLMC in the case of the Lotka-Volterra and M/G/1 queue models, we find that the unbiased method can greatly benefit from variance reduction. In the case of Lotka-Volterra, when comparing TGRR and GRR methods, we conclude that the gradient information $\Delta\rho_{\phi,\ell}$ where $\ell\geq\overline{m}$ not only does not contribute to the overall performance but also make it worse when measured with C2ST. As suggested by Theorem 4.2, this could be due to the domination of variance, where the effect of reducing variance is superior to reducing bias.
In conclusion, in the case of SNPE where the complex model of $q_{F(x,\phi)}(\theta)$ is used for density estimation, variance tends to dominate the bias as it is suggested by Theorem 4.2, indicating that excessive variance could seriously affect the training process of the density estimator. Instead of seeking unbiasedness, one should try to strike a balance between bias, average computational cost, and variance of the gradient in this case.
It is worth noting that in the case of the Two-moon problem, RU-MLMC method delivers comparable performance to the other two methods with lower computational cost and time requirements. Rather than employing any of these methods indiscriminately, it is crucial to undertake a comprehensive and thoughtful analysis of the specific problem at hand.
In Table 2, we also compare TGRR-MLMC with sequential neural score estimation (SNSE) proposed by [54], which is a state-of-the-art likelihood-free inference. We find that TGRR-MLMC demonstrates consistent improvement over SNSE across three baseline tasks. As shown in Table 3, TGRR-MLMC takes approximately 1.2–1.5 fold computational time of SNSE. However, SNSE consumes up to 60GB of GPU memory for a batch size of 128, while our method requires only 5GB for a batch size of 500, which demonstrates a substantial advantage in memory efficiency. For the hyperparameter of SNSE, we mainly follow the default setting from [13] and change the maximum iteration to 1,000 and samples per round to 1,000 to align with our settings.
To investigate the performance of the proposed MLMC method scaling with higher dimensional problems, we infer the single-compartment Hodgkin-Huxley type neuron model [49, 23] with 8 parameters of interest. The results are deferred to Appendix E.
Table 2: C2ST of different methods after 10,000 simulations. The values represent the mean and standard deviation.
Task Two-moon Lotka-Volterra M/G/1
TGRR-MLMC $0.5407\pm 0.0315$ $0.7472\pm 0.0599$ $0.6402\pm 0.0210$
SNSE $0.5598\pm 0.0673$ $0.9153\pm 0.0280$ $0.8394\pm 0.0116$
Table 3: Comparison of computational time costs (in minutes). The values represent the mean and standard deviation.
Methods/Tasks Two-moon Lotka-Volterra M/G/1
TGRR-MLMC $299.35\pm 28.96$ $300.10\pm 26.88$ $294.65\pm 18.67$
SNSE $194.76\pm 12.76$ $217.64\pm 18.20$ $243.80\pm 17.28$
@@ 6 Concluding remarks
The atomic APT method is a heuristic nested estimation algorithm. Given that its algorithm design is not friendly to theoretical analysis, this paper first proposes a nested APT method with similar performance, thus enabling a series of convergence analyses to quantify the impact of inner and outer sample sizes on the computational complexity. Given the biased nature of the nested APT method, this paper conducts a comprehensive exploration based on a series of unbiased MLMC to eliminate the bias in the nested APT method. This paper also proposes several theorems to quantify the bias, variance, and average computational cost of the loss function and its gradient estimators. Additionally, since the optimization method based on SGD is used to update the neural network parameters, it is well-known that unbiased methods suffer from excessively high variance in gradient estimation. To address this issue, this paper proposes using several variants of the truncated MLMC method to reduce the variance. A series of numerical experiments verify the effectiveness of these variants.
We acknowledged some limitations regarding the proposed MLMC methods. The methods may suffer from excessive variance accompanied by high dimensional problems and complex neural network structures, which lead to inferior performances in some extent. A possible way to reduce the variances of MLMC estimators is to use quasi-Monte Carlo (QMC) methods, which use low discrepancy (more balanced) points to generate samples from the proposal. QMC-based quadrature can yield a faster convergence rate than plain Monte Carlo, and its application in nested MLMC methods have already been investigated in [29]. Secondly, MLMC takes more computational time than atomic APT due to the gradient calculation at multiple levels. How to speed up the algorithm of MLMC is left for future work.
Appendix A Proof for Theorem 3.2
We first recall a useful lemma for bounding moments and tail probability of iid sample mean, which was used in [18].
Lemma A.1.
Let $X$ be a real-valued random variable with mean zero, and let $\overline{X}_{N}$ be an average of $N$ iid samples of $X$. If for $\mathbb{E}[|X|^{u}]<\infty$ for $u\geq 2$, there exists a constant $C_{u}>0$ depending only on $u$ such that
for any $c>0$.
Proof of Theorem 3.2.
The proof follows [22]. We first note that
$M_{\max}$ $:=\underset{x,\phi,\theta}{\sup}\|\nabla_{\phi}\log g_{\phi}(x,\theta)\|_{\infty}=\underset{x,\phi,\theta}{\sup}\frac{\|\nabla_{\phi}g_{\phi}(x,\theta)\|_{\infty}}{g_{\phi}(x,\theta)}<\infty,$ (35)
@@
$H(u)$ $:=\underset{\phi\in\Phi}{\sup}\ \mathbb{E}\left[\left(\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right)^{u}\right]<\infty,\ \forall u\in(0,s].$
Define the event
where $S^{(i)}_{\ell-1}:=g^{(i)}_{\phi,M_{\ell-1}}(x)/Z(x,\phi)-1$ with $i\in\{a,b\}$, which is a sample mean of $M_{\ell-1}$ iid copies of $g_{\phi}(x,\theta)/Z(x,\phi)-1$ with zero mean. We also define
$S_{\ell}:=\frac{g_{\phi,M_{\ell}}(x)}{Z(x,\phi)}-1=\frac{S^{(a)}_{\ell-1}+S^{(b)}_{\ell-1}}{2}.$
Let $\mathcal{A}^{c}$ be the complement of the event $\mathcal{A}$. Thanks to the following decomposition
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\right]=\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\mathbf{1}_{\mathcal{A}}\right]+\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\mathbf{1}_{\mathcal{A}^{c}}\right],$ (36)
it suffices to develop a proper upper bound for the two terms.
[…]
$\frac{\sum_{j=1}^{d}a_{j}}{\sum_{j=1}^{d}b_{j}}\leq\underset{i=1,\dots,d}{\max}\left\{\frac{a_{i}}{b_{i}}\right\}.$ (37)
Using (37), we have
$\left\|\nabla_{\phi}\log\hat{Z}_{M}(x,\phi)\right\|^{2}_{2}$ $=\left\|\frac{\sum_{j=1}^{M}\nabla_{\phi}g_{\phi}(x,\theta_{j}^{\prime})}{\sum_{j=1}^{M}g_{\phi}(x,\theta_{j}^{\prime})}\right\|^{2}_{2}=\sum_{i=1}^{d}\left(\frac{\sum_{j=1}^{M}|\partial_{\phi_{i}}g_{\phi}(x,\theta_{j}^{\prime})|}{\sum_{j=1}^{M}g_{\phi}(x,\theta_{j}^{\prime})}\right)^{2}$
$\leq d\left\|\nabla_{\phi}\log g_{\phi}(x,\theta)\right\|_{\infty}^{2}\leq dM_{\max}^{2},$
where $\phi=(\phi_{1},\dots,\phi_{d})\in\mathbb{R}^{d}$. As a result,
$\left\|\rho_{\phi,M}\right\|^{2}_{2}$ $=\left\|-\nabla_{\phi}\log g_{\phi}(x,\theta)+\nabla_{\phi}\log\hat{Z}_{M}(x,\phi)\right\|^{2}_{2}$
$\leq 2\left\|\nabla_{\phi}\log g_{\phi}(x,\theta)\right\|^{2}_{2}+2\left\|\nabla_{\phi}\log\hat{Z}_{M}(x,\phi)\right\|^{2}_{2}\leq 4dM_{\max}^{2}.$
This gives
It then follows
$\leq 3\left\|\rho_{\phi,M_{\ell}}\right\|_{2}^{2}+\frac{3}{4}\left\|\rho^{(a)}_{\phi,M_{\ell-1}}\right\|_{2}^{2}+\frac{3}{4}\left\|\rho^{(b)}_{\phi,M_{\ell-1}}\right\|_{2}^{2}\leq 18dM^{2}_{\max}.$ (38)
For any $u\in[2,s]$, applying Lemma A.1 gives
$\mathbb{P}\left[\mathcal{A}\right]$ $\leq\mathbb{P}\left[\left|S^{(a)}_{\ell-1}\right|>\frac{1}{2}\right]+\mathbb{P}\left[\left|S^{(b)}_{\ell-1}\right|>\frac{1}{2}\right]$
$\leq\frac{2C_{u}}{2^{-u}M_{\ell-1}^{u/2}}\mathbb{E}[|g_{\phi}(x,\theta)/Z(x,\phi)-1|^{u}]$
$\leq\frac{2^{3u/2+1}C_{u}}{M_{\ell}^{u/2}}(\mathbb{E}[(g_{\phi}(x,\theta)/Z(x,\phi))^{u}]+1)$
$\leq\frac{2^{3u/2+1}C_{u}}{M_{\ell}^{u/2}}(H(u)+1).$ (39)
Then for the first term in the right hand side (RHS) of (36), we have
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\mathbf{1}_{\mathcal{A}}\right]\leq 18dM^{2}_{\max}\mathbb{P}\left[\mathcal{A}\right]=\mathcal{O}\left(M_{\ell}^{-s/2}\right).$ (40)
For the second term $\mathbb{E}[\|\Delta\rho_{\phi,\ell}\|_{2}^{2}\mathbf{1}_{\mathcal{A}^{c}}]$ in the RHS of (36), we utilize the antithetic property to obtain the following identity
$\Delta\rho_{\phi,\ell}$ $=\frac{1}{2}\tilde{F}_{\ell-1}^{(a)}S_{\ell-1}^{(a)}+\frac{1}{2}\tilde{F}_{\ell-1}^{(b)}S_{\ell-1}^{(b)}-\tilde{F}_{\ell}S_{\ell}$
$\quad+\frac{1}{2}\frac{\nabla_{\phi}Z(x,\phi)}{g^{(a)}_{\phi,M_{\ell-1}}(x)}\left(S_{\ell-1}^{(a)}\right)^{2}+\frac{1}{2}\frac{\nabla_{\phi}Z(x,\phi)}{g^{(b)}_{\phi,M_{\ell-1}}(x)}\left(S_{\ell-1}^{(b)}\right)^{2}-\frac{\nabla_{\phi}Z(x,\phi)}{g_{\phi,M_{\ell}}(x)}\left(S_{\ell}\right)^{2},$ (41)